OpenStudy (anonymous):

WILL GIVE MEDAL A customer at an online movie rental company pays a total of $25 after 1 month, a total of $30 after 2 months, and a total of $35 after 3 months. The total amount paid is represented by T and the number of months by m. Which equation models the total payment, T, after m months? A. T = 10m + 15 B. T = 5m + 15 C. T = 5m + 20 D. T = 10m

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@studyislandhelp

OpenStudy (raffle_snaffle):

@Astrophysics

OpenStudy (anonymous):

C is the answer because when you substitute the totals ($25, $30, $35) for T and solve each equation you will get the corresponding months (m) of {1, 2, 3} respectfully, which equally model total payments, T, after each month.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

can u help me with this one @VAStaten Almonds sell for $4.50 per pound, and green beans for $3.75 per pound. Paul bought some almonds and some green beans. The total weight was 3.5 pounds and cost $9.75. Let p represent the number of pounds of almonds. Which equation represents the situation described? A. 3.75p + 4.5(p – 3.5) = 9.75 B. 4.5p + 3.75(p – 3.5) = 9.75 C. 4.5p + 3.75(3.5 – p) = 9.75 D. 3.75p + 9.75(3.5 – p) = 4.5

OpenStudy (anonymous):

you there @VAStaten

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

do u know the anwswer

OpenStudy (anonymous):

C

OpenStudy (anonymous):

are u sure

OpenStudy (anonymous):

last one Glenn works as an office clerk and his office job pays him $12 per hour. Sheila works as a cashier and her job pays her $13.50 per hour. Together, they worked 55 hours and earned a total of $705. How many hours did Glenn work in his job? A. 20 B. 25 C. 30 D. 35

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

B

OpenStudy (anonymous):

u sure

OpenStudy (anonymous):

this is my last time taking it

OpenStudy (anonymous):

Glenn worked 25 hours which is 25 x $12 =$300 the other person worked 30 hours , which is 30 x $13.50 = $405 thus 300 + 405 = 705 and 55 hours total

OpenStudy (anonymous):

ok i will tell if i past

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

ok i will tell if i past

OpenStudy (anonymous):

i past thank you soooo much

OpenStudy (anonymous):

you are awesome

OpenStudy (anonymous):